Solve the equation: the hypotunus is 13. b= x+7 and A=x
Answer by checkley75(3666)

a^2+b^2=c^2
x^2+(x+1)^2=13^2
x^2+x^2+2x+1=169
2x^2+2x+1-169=0
2x^2+2x-168=0
using the quadratic equation we get:
x=(-2+-sqrt[2^2-4*2*-168])/2*2
x=(-2+-sqrt[4+1344])/4
x=(-2+-sqrt1348)/4
x=(-2+-36.66)/4
x=(-2+36.66)/4
x=34.66/4
x=8.67 answer
x=(-2-36.66)/4
x=-38.66/4
x=-9.67 non answer.